OpenStudy (anonymous):

Find the equation of the line that has a slope of 7 and passes through the point (-8,3)

OpenStudy (anonymous):

Have you done it yet? Do it on the graph and tell me.

OpenStudy (anonymous):

If you have an answer choice please give it.

OpenStudy (anonymous):

i typed the wrong question

OpenStudy (anonymous):

can you help me real quick on this one

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

Find the equation of the line that has a slope of -4 and a y intercept of 5

OpenStudy (anonymous):

Do you have an answer choice, a picture?

OpenStudy (anonymous):

no all i have is examples of how to do it and they dont make sense

OpenStudy (anonymous):

Usually, they should put an answer choice it would help a lot.

OpenStudy (anonymous):

but i have to find the equation of the line

OpenStudy (anonymous):

Hmm.. I need some hints to help me solve it.

OpenStudy (anonymous):

can i give you the examples they gave me

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

Find the equation of the line use either point: (3,5) remember the form: y-y1=m(x-x1) Substitute: y-5 = 6.5 (x-(-3)) y-5=6.5 (x+3) y-5 = 6.5x +19.5 To write in Y form: (continue to simplify algebraically) y=6.5x + 19.5 +5 y=6.5x+24.5 answer= y=6.5x+24.5

OpenStudy (anonymous):

This is easy.

OpenStudy (anonymous):

y=-4+5

OpenStudy (anonymous):

Find the equation of the line that has a slope of -4 and a y intercept of 5

OpenStudy (anonymous):

thats how you do it?

OpenStudy (anonymous):

Yes because it's finding the equation. Just remember, this is a slope-intercept form: y = mx + b. mx is the slope b is the y-intercept.

OpenStudy (anonymous):

so -4+5

OpenStudy (anonymous):

Yes, put it the way I showed you it. y=-4+5

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

You're welcome. Don't forget to give me a medal and fan me if you have another question.

OpenStudy (anonymous):

did i do that right